40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9251     Entry time: Thu Oct 17 18:29:28 2013
Author: Koji 
Type: Update 
Category: LSC 
Subject: PRMI+2arm attempt (not really yet) 

While Manasa, Jenne, and Masayuki are working on the preparing the interferometer, I write the elog for them.

- 6PM-ish: X and Y arms were was locked. They were aligned with ASS.

- PRMI was locked. The PRM was aligned with ASS.

- Jenne went into the lab and aligned the PRM ASC QPD.

- Jenne also aligned all of the oplev spots except for the SRM.

- 6:40PM Then, Manasa and Masayuki checked the out-of-loop stability of the arms.
The X and Y arms have the rms of 2.2kHz and 600Hz, respectively.
The X arm is significantly worse than the Y arm.

Masayuki saved the plot somewhere in his directory.

- 7:20PM X beat: 41.2MHz, Y beat: 14.8MHz

- 7:22PM PRMI locked POP110 115-120 

- 7:30PM Lost lock of everything. Start over. Taking the arm alignment.

- 7:45PM start the 2nd trial. PRMI+one arm ready.

- 8:00PM explosion! Lost lock.

- 8:30PM The Xarm ALS is not stable anymore. It loses the control in ~10sec.
We are investigating the out-of-loop stability of the Yarm ALS.
(i.e. Look at the beat note error signal while locking the Yarm with the IR PDH)

ELOG V3.1.3-